OpenStudy (anonymous):

if sin theta = 4/5 and is in quadrant 2, what is the value of tan theta

OpenStudy (anonymous):

\[\huge \sin \theta = \frac{4}{5}\] we know that: \[\huge \sin^2 \theta + \cos^2 \theta = 1\] \[\huge \rightarrow \cos^2 \theta = 1-\sin^2 \theta= 1-(\frac{4}{5})^2\] \[\huge \rightarrow \cos^2 \theta = 1-\frac{16}{25}= \frac{25-16}{25}= \frac{9}{25}\] \[\huge \rightarrow \cos^2 \theta = \frac{9}{25}\] \[\huge \rightarrow \cos \theta = \pm \sqrt{ \frac{9}{25}}\] \[\huge \rightarrow \cos \theta = \pm \frac{3}{5}\] Since theta lies in the second quadrant, hence \[\huge \rightarrow \cos \theta = - \frac{3}{5}\] Now we know that: \[\huge \tan \theta = \frac{\sin \theta }{ \cos \theta}\] \[\huge \rightarrow \tan \theta = \frac{\frac{4}{5} }{\frac{-3}{5}}\] \[\huge \rightarrow \tan \theta = \frac{4}{5} \times \frac{5}{-3}\] \[\huge \rightarrow \tan \theta =\frac{4}{-3}\] \[\huge \rightarrow \tan \theta =-\frac{4}{3}\] @TSV